OpenStudy (anonymous):

Find the angle between each of the following pairs of vectors A⃗ =Axi^+Ayj^ and B⃗ =Bxi^+Byj^. Ax1= -3.20, Ay1= 4.40; Bx1= 2.70, By1= -3.10 i know this is dot prodoct but the calculater just keep saying error

OpenStudy (zzr0ck3r):

what is the error?

OpenStudy (zzr0ck3r):

\(a\cdot b = -3*2.70+4.40*-3.10 = -21.74\) so \(\theta=\cos^{-1}(\frac{-21.74}{\sqrt{3^2+4.4^2}*\sqrt{2.7^2+3.1^2}})=172^{\circ}\)

OpenStudy (zzr0ck3r):

boo at post and leave...